OpenStudy (anonymous):

can someone please help. (x-(1/y))/(x/y) = ?

OpenStudy (anonymous):

hello :)

OpenStudy (anonymous):

groges help

OpenStudy (anonymous):

(xy-1)/x ?

OpenStudy (anonymous):

\[\frac{ x-\frac{ 1 }{ y } }{ \frac{ x }{ y } }\] Is this your problem?

OpenStudy (anonymous):

@whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 @KawasumiKimiko @Luigi0210 @Loujoelou @love_jessika15 @lkkk @Callisto @countonme123 @ChrisJP @nincompoop @mukushla @marylou004 @mlb004 @MissHN182 @thefon19 @ryanjnfish @ladycomical @DDCamp @ryanjnfish @jd1028 @phantomofthemathwork @jtrich13 @ajprincess @PaulaCamacho @laxman03 @beccabee77 @babyslapmafro @whpalmer4 @jtrich13 @resident2013 @sumi29 @R00nnYBraiNsbiG @tfuller @Frostbite @SnuggieLad @raque766 @physicsboffin @Falco276 @e.cociuba @blurbendy @kausarsalley @sakigirl @RANE @Danny_Boy @SomeGirl1999 @Jamierox4ev3r @iforgot @lilkittykat36 @n14r96 @ChanceBBrick @knowel @meldo16 @Helpmeplease99 @Manda_luvsu @baseballer20142 @jhalt @RIUKUS @harpelletjain @Math81 @Frank @WildKat @maurimath @shamim @AliaKnight @MayMay_69 @amyamanda @victoria650 @Social @Whatever1 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @Zoe_Bailey @RIUKUS @henendez @shaneetta @niah2411 @harpelletjain @Math81 @WildKat @eidan @Hassan07 @Asuman @Frank @bani2100 @turtlebuddy @maurimath @lizzy89 @dborder @peush @EricaaaaB @yehonatan @tori45 @Ojasmin @Markbruce @april115 @RH @shamim @Ryaan @emilyhaddad @Jamierox4ev3r @RANE @NoelGreco @shkrina @lakenclark @Fazeelayaz @sarahusher @JazzyFrizo16 @music101 @HaperFink22 @whydoihavetosignup1 @domu @ashleyjade @Gikundab @Will.Neko @QuinterrianW_23 @tillman--- @baseballer20142 @Manda_luvsu @Charollete @amyamanda @harpreetsk @Whatever1 @HaydenBurden1996 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @shaneetta @DLS @NoelGreco @emilyhaddad @marylou004 @JazzyFrizo16 @music101 @HaperFink22 @Courage. @KingAditya @bradely @Gikundab @Will.Neko @QuinterrianW_23 @mcordray @_anastasia0 @ozhobbits @gabigurl @musicalcat98 @Obviously717 @annemoreno @glamourariss @jbsharpe @WildKat @turtlebuddy @turtlebuddy @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Ojasmin @Markbruce @confusao @sircrashalot @iizabell @tori45 @yehonatan @peush @maurimath @jbsharpe @glamourariss @annemoreno @Obviously717 @mcordray @_anastasia0 @AppylPye @gabigurl @Imtiaz7 @fringe @khadeeja @jhalt @itsmeeQUENNIE @KingAditya @sambhav__jain @whydoihavetosignup1 @mlb004 @Yasminelove1 @caozeyuan @marylou004 @genius12 @agent0smith @AnElephant @lakenclark @JazzyFrizo16 @HaperFink22 @music101 @danielle02 @HaydenBurden1996 @xxAshxx @Obviously717 @Zoe-bailey @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Fabien1 @Ojasmin @sircrashalot @ankit042 @caozeyuan @rperez36 @freddy_eighty7 @Charollete @freddy_eighty7 @toolCoolChris @agent0smith @genius12 @uri @walac @Luigi0210 @kelliegirl33 @nincompoop @mathmate @stgreen @henpen @binary3i @ChrisJP @zpupster @ankit042 @caozeyuan @betterwiththelightsoff99 @Yasminelove1 @HaperFink22 @glitterythings @whydoihavetosignup1 @Charollete @bradely @ashleyjade @Alexis2332 @Will.Neko @walac @Gikundab @Imtiaz7 @Dianatorres @Alyssa_May_De_Jesus @baseballer20142 @victoria650 @lakenclark @zpupster @mathmate @binary3i @austinL @Loser66 @GoldPhenoix @e.cociuba @DLS @razor99 @Mashy @Machida @RANE @Goten77 @lncognlto @Fifciol @ivancsc1996 @kenbaby @kryton1212 @ankit042 @Henry.Lister @JazzyFrizo16 @betterwiththelightsoff99 @mlb004 @ivano_logy @xxkerstie.bbyxx @MayMay_69 @KawasumiKimiko @itsmeeQUENNIE @Shadly @Albany_Goon @10thgrader @xxmarcusxx1994 @jessy1289 @baseballer20142 @groges @pdd21 @claringleek @Rowa @daniellecrook @imcool @harpreetsk @angelj @JJB @OpenSessame @xseahaven @Whatever1 @harliedanielfrye @borimexz @Jdubose2 @BrianaShamwell @freddy_eighty7 @Horush @snewton29 @Fabien1 @Peter_szlapak @rmpjingwei @mtalampas21 @andriani @mygumshurt @lamel1993 @LittleAnnn @psychomath @yehonatan @heights @JupitersBrain @slayer00647 @2old4math @munastar @jmensal @TanteAnna @Matt-Teepell @DJova @quietkat @bcrisco @larlar77 @Markevia @utvet @felixt03 @9995950976 @natyxz @lftyoutofthepen @kiah44 @Henry.Lister @helder_edwin @Loser66 @inkyvoyd @mathmate @hartnn @dmezzullo @abb0t @amistre64

OpenStudy (anonymous):

how did you get the groges? and yes it is agent

OpenStudy (anonymous):

First, in the numerator what could we do to combine everything? Second, the denominator has a fraction. Remember that dividing by a fraction is the same as multiplying by it's reciprocal. Does that make sense?

OpenStudy (anonymous):

yes. so far so good

OpenStudy (anonymous):

That's all you gotta do. :D Some terms my 'cancel' out.

OpenStudy (anonymous):

ok stupid question then. when you multiply the top by the flipped denominator, does it apply to both terms on the top?

OpenStudy (anonymous):

\[\frac{ x-\frac{ 1 }{ y } }{ \frac{ x }{ y } }=(x-\frac{ 1 }{ y })\frac{ y }{ x } = \frac{ y }{ y }(x-\frac{ 1 }{ y })(\frac{ y }{ x })=\frac{ yx-y }{ y }\cdot \frac{ y }{ x }\] I think you could do the rest from here. I just wanted to make sure the concept was demonstrated. :D

OpenStudy (anonymous):

Not a stupid question. Yes it would.

OpenStudy (anonymous):

could you finish it? like do you multiply next?

OpenStudy (anonymous):

Lets see if you can get it finished. First, did all of my steps make sense to you?

OpenStudy (anonymous):

yes, multiplying by y/y to give a common denominator.

OpenStudy (anonymous):

Great. Honestly I think you can do this on your own. Just got to simply it by possibly canceling out similar terms on top and bottom.

OpenStudy (anonymous):

I got it now! I was having the biggest brain fart ever!!! thanks so much!

OpenStudy (anonymous):

No problem man. Glad to see you got it all on your own! :D